BexleyBrooks:

Need help with some math lol. Alistair has a cube-shaped box that has a volume of 216 cubic inches. 1. How would you write the edge length of Alistair's box using a cube root? 2. Fill in the x's 3 divided by 216=3 divided by 6 times x times x=3 divided by x to the third power = x

silvernight269:

Ok so i can help you with part one. I do bot understand what part 2 wants.

BexleyBrooks:

@silvernight269 wrote:
Ok so i can help you with part one. I do bot understand what part 2 wants.
so for number 2 it wants me to fill in the x's of the problem

silvernight269:

Ok so for part one: lets say we wanted to find the area of a square. There are 2 ways we can write for the formula. One would be LxW=A. But another way would be L^2=A because the length and width equals the same. 2x2 can be written as 2^2. 3x3 —> 3^2 or 2x2x2 can be 2^3. Do you know where Im going with this?

silvernight269:

@bexleybrooks wrote:
@silvernight269 wrote:
Ok so i can help you with part one. I do bot understand what part 2 wants.
so for number 2 it wants me to fill in the x's of the problem
No literally all i see is: 216/3=6XxX/3=X^3/3=x

BexleyBrooks:

@silvernight269 wrote:
Ok so for part one: lets say we wanted to find the area of a square. There are 2 ways we can write for the formula. One would be LxW=A. But another way would be L^2=A because the length and width equals the same. 2x2 can be written as 2^2. 3x3 —> 3^2 or 2x2x2 can be 2^3. Do you know where Im going with this?
I do see where your going with it

silvernight269:

@bexleybrooks wrote:
@silvernight269 wrote:
Ok so for part one: lets say we wanted to find the area of a square. There are 2 ways we can write for the formula. One would be LxW=A. But another way would be L^2=A because the length and width equals the same. 2x2 can be written as 2^2. 3x3 —> 3^2 or 2x2x2 can be 2^3. Do you know where Im going with this?
I do see where your going with it
Do you think you can find the answer for part 1 then?

BexleyBrooks:

@silvernight269 wrote:
@bexleybrooks wrote:
@silvernight269 wrote:
Ok so for part one: lets say we wanted to find the area of a square. There are 2 ways we can write for the formula. One would be LxW=A. But another way would be L^2=A because the length and width equals the same. 2x2 can be written as 2^2. 3x3 —> 3^2 or 2x2x2 can be 2^3. Do you know where Im going with this?
I do see where your going with it
Do you think you can find the answer for part 1 then?
yes

BexleyBrooks:

@silvernight269 wrote:
@bexleybrooks wrote:
@silvernight269 wrote:
Ok so i can help you with part one. I do bot understand what part 2 wants.
so for number 2 it wants me to fill in the x's of the problem
No literally all i see is: 216/3=6XxX/3=X^3/3=x
\[\int\limits_{3}^{}216=\int\limits_{3}^{}6\times x \times x=\int\limits_{3}^{}0 ^{3}=x\] It was me to fill in all the x's

silvernight269:

@bexleybrooks wrote:
@silvernight269 wrote:
@bexleybrooks wrote:
@silvernight269 wrote:
Ok so for part one: lets say we wanted to find the area of a square. There are 2 ways we can write for the formula. One would be LxW=A. But another way would be L^2=A because the length and width equals the same. 2x2 can be written as 2^2. 3x3 —> 3^2 or 2x2x2 can be 2^3. Do you know where Im going with this?
I do see where your going with it
Do you think you can find the answer for part 1 then?
yes
Great

silvernight269:

@bexleybrooks wrote:
@silvernight269 wrote:
@bexleybrooks wrote:
@silvernight269 wrote:
Ok so i can help you with part one. I do bot understand what part 2 wants.
so for number 2 it wants me to fill in the x's of the problem
No literally all i see is: 216/3=6XxX/3=X^3/3=x
\[\int\limits_{3}^{}216=\int\limits_{3}^{}6\times x \times x=\int\limits_{3}^{}0 ^{3}=x\] It was me to fill in all the x's
Uhhh those fs ain’t for algebra 😭 those are for derivatives

silvernight269:

Ain’t no way am i going back to calculus books 🫡 sorry 🙇🏻‍♂️

BexleyBrooks:

@silvernight269 wrote:
Ain’t no way am i going back to calculus books 🫡 sorry 🙇🏻‍♂️
It's totally fine haha. They made it super hard and I don't understand why they put those 2 different types of math tg

OLIVER69:

If question is answered please close this question, so you are able to ask more questions. If not then I may help with the next part if I know what I need to explain to you next...